# HERACLES++: a multi-dimensional Eulerian code for exascale computing

### v1

@@ 4.3 Three-dimensional off-centered Sedov blast wave test
The test of the Sedov blast wave (Sedov, 1946) is useful to describe the energy-conserving phase of a SN remnant. We perform this test using a grid in spherical geometry in order to test the precision of our numerical scheme for a non-Cartesian grid. The problem consists of a stationary background medium with an uniform density $\rho_{0}=1$ and a low pressure $P_{0}=0.01$. For our test, the domain is a spherical wedge with $r\in[0.1,1.3]$ and $(\theta,\phi)\in[\pi/4,3\pi/4]$. The blast is created with a peak of total energy $E_{1}=1$ at position $\vec{r_{0}}=(r_{0},\theta_{0},\phi_{0})=(0.7,\pi/2,\pi/2)$. The boundary conditions are transmissive in every directions, although the simulation is stopped before the blast wave reaches any boundary. The spherical polar grid counts 256 zones in all three directions.
The distance between the shock and $\vec{r_{0}}$ is given analytically by:
$d_{\rm sh}(t)=\xi_{0}\Big{(}\frac{E_{1}\,t^{2}}{\rho_{0}}\Big{)}^{1/5}$ (25)
@@
Figure 5 shows a meridional slice of the density at $t=0.08$ together with the analytical solution given by Eq. 25. Other slices through the 3-D domain yield a similar pattern. Hence, the simulation is in good agreement with the analytical prediction, apart from a small offset in the direction where the shock is aligned with the grid. In such locations, the numerical diffusion is lower, causing a notorious problem for numerical schemes (Elling, 2009).
Figure 5: Meridional slice of the density for the 3-D off-centered Sedov blast wave test. The time is 0.08 and the cut is done along $\phi=\pi/2$. The analytical prediction given by Eq. 25 is shown as a black line. (See §4.3 for discussion.)
@@
Figure 6: Initial conditions for our HERACLES++ simulations. We show the density (top left), temperature (top right), velocity (bottom left) and the mass fraction of H, He, O, Si and 56Ni, as a function of Lagrangian mass. The time is 500 s after the explosion trigger.
@@ 5.1 Preamble
We now turn to the more astrophysics-oriented context of a spherical explosion occurring in a RSG star. This initial work is used here for the benchmarking of the code but it fits within a larger project of simulating the various instabilities taking place in the extended envelopes of RSG stars before, during, and after their explosion. These so-called Type II SNe are the most frequently encountered type of SNe from massive star explosions (see, e.g., Perley et al. 2020). Furthermore, detailed NLTE radiative-transfer calculations have demonstrated that both small and large-scale structures of SN ejecta have a significant impact on SN light curves and spectra (Dessart et al., 2018; Dessart & Audit, 2019; Ergon & Fransson, 2022) so progress in that sector is highly needed in order to improve the physical consistency of radiative-transfer and radiation-hydrodynamics simulations and the robustness of their inferences.
For this work, we used the initial ejecta conditions of Dessart et al. (2024), which are based on a subset of evolutionary calculations for binary massive stars performed by Ercolino et al. (2024). Binarity, which occurs with a high frequency in massive stars (Sana et al., 2012), can lead to a wide range of H-rich envelope mass $M$(H-env) in the preSN star because of mass transfer through Roche-lobe overflow. However, as long as a residual H-rich envelope remains, the He-core mass as well as the surface radii are largely unaffected by this mass loss, so that these objects should explode with similar properties (explosion energy and nucleosynthesis), merely modulated by the differing $M$(H-env). Dessart et al. (2024) indeed showed that through variations in $M$(H-env) alone, the full diversity of Type II SN light curves may be reproduced. Chemical mixing and clumping may, however, differ widely between these ejecta due to the differences in their progenitor-envelope properties, with strong implications for their observables. Investigating this diversity will be the subject of a forthcoming study.
@@
In practice, we started our simulations from the output of the explosion simulation carried out with the 1-D, Lagrangian radiation-hydrodynamics code V1D (Livne, 1993; Dessart et al., 2010a, b), at a time of 500 s relative to the onset of the explosion trigger. The total energy injected in the star for the explosion was designed to yield a kinetic energy at infinity of $1.1\times 10^{51}$ erg so by this time of 500 s, the shock has already reached the outer layers of the He core [^3] For details about the preSN evolution and explosion simulation, we refer the reader to Dessart et al. (2024) and Ercolino et al. (2024).. Using HERACLES++, we first evolved this structure in 1-D (Section 5.3) and then in 3-D (Section 5.5) in order to study the development of Rayleigh-Taylor instabilities and their impact on chemical mixing and clumping.
Figure 7: Variation of the radial grid spacing $dr$ as the function of radius $r$. The grid used in the simulation with HERACLES++ is regular until 10 $R_{\odot}$ and logarithmic beyond.
Figure 8: Ejecta properties for our RSG star explosion model as computed by HERACLES++ (solid) and V1D (dashed). For representative epochs between 2.8 hr and 1.6 d, we show the density (top row), the temperature (middle row) and the velocity (bottom row) versus Lagrangian mass (left column) and radius (right column). (See §5.3 for discussion.)
@@ 5.2 Initial conditions and numerical setup
The ejecta properties used as initial conditions for the HERACLES++ simulations are shown in Fig. 6. The shock is located at $\sim$ 4.1 $M_{\odot}$ with a peak velocity around 12 000 km s-1. Beyond the shock the velocity is essentially zero and $\rho$ and $T$ slowly decrease until the progenitor surface at 10.84 $M_{\odot}$. Figure 6 also shows the mass fraction of the main elements (i.e., H, He, O, Si, and ${}^{56}\rm{N}$i). The position of each layer in the star reflects previous episodes of nuclear burning or mixing, as well as the ashes of explosive nucleosynthesis in the innermost ejecta layers. The extended H-rich envelope is essentially unchanged from its original composition (i.e., corresponding here to solar metallicity), whereas the He-core layers exhibit a stratification of ashes from H, He, C, O, Ne, and Si burning. The explosion energy at 500 s is primarily stored as kinetic energy, with around $9\times 10^{50}$ erg.
@@
Since our interest is to investigate the impact of fluid instabilities on the distribution of elements, we introduced five passive scalars to track the material originally present in the H-rich envelope and in the He-, O-, Si-, or ${}^{56}\rm{N}$i-rich shells. In the following sections, when referring to the distribution of O, we really mean the distribution of the material in the O-rich shell (in practice, this includes the O/Si, O/Ne/Mg and the O/C shells).
@@
Figure 9: Evolution of the total energy $E_{\rm tot}$, the kinetic energy $E_{\rm kin}$, the radiative energy $E_{\rm rad}$ and the internal (gas) energy $E_{\rm gas}$ during our HERACLES++ simulation from 500 s until shock breakout. At shock breakout, $E_{\rm kin}\sim E_{\rm rad}$.
@@
At the start of the HERACLES++ simulation, we remapped the V1D input (which is on a Lagrangian grid) onto the HERACLES++ grid, which we chose to be a hybrid linear/log grid. It has a relatively high resolution out to the shock location (in order to preserve the initial sharpness of the shock and the total initial energy of the explosion) and a decreasing resolution beyond. In Section 5.3, we constructed a regular radial grid with 10000 points to produce a grid that presents no major resolution offset with the V1D grid, which is Lagrangian and thus in mass space (i.e., the goal in this test is to confront the numerical solution of the hydrodynamical equations between the two codes, rather than assess the influence of resolution or numerical diffusion etc). In Section 5.5, we created a radial grid with a total of 1024 points, with 200 points for the regular part out to 10 $R_{\odot}$ (corresponding to a Lagrangian mass of $\sim$ 4.2 $M_{\odot}$). Figure 7 illustrates the variation of the radial increment $dr$ with radius. In Appendix A, we consider additional radial resolutions (i.e., $n_{r}$ of 256, 512, and 768) as part of a resolution study.
@@
We performed hydrodynamical simulations with HERACLES++, using spherical coordinates, first in 1D (Section 5.3) and then in 3D (Section 5.5). We used the perfect gas+radiation EoS (Eq. 6b; we used $\gamma=5/3$ and $\mu=0.62$ for the ideal-gas part) and introduced ${}^{56}\rm{N}$i-decay heating (Eq. 8; this was done for consistency with V1D but it plays little role here since we stopped the simulation at about 1.5 d, which is only a quarter of the ${}^{56}\rm{N}$i lifetime). A point mass gravity (Eq. 4) was used to account for the presence of the compact remnant. We used a routine to track precisely the internal energy (§2.5). The boundary conditions were reflexive in the inner radius and transmissive at the outer radius. The slope limiter used was Minmod and the Riemann solver was HLLC. For consistency, this general setup was the same for all similar simulations.
@@
Figure 10: Estimated growth rate of the R-T instability in our RSG explosion simulation, shown at multiple epochs and versus Lagrangian mass. (See §5.4 for discussion.)
@@ 5.3 Results in spherical symmetry and comparison with V1D
In this section, we first present a comparison of the results between V1D and HERACLES++, and thus in spherical symmetry. Figure 8 shows the evolution of $\rho$, $T$ and $u$ over time as the shock propagates through the progenitor envelope and until its emergence at $R_{\star}$ at $\sim$ 1.6 d. [^4] For this figure and to ensure the Eulerian grid is as fine in the inner regions as the Lagrangian grid in V1D, the HERACLES++ simulation employs 10 000 radial zones. We show the variations of these quantities both as a function of Lagrangian mass and radius. There is a good agreement between the results of the two codes.
A pair of reverse/forward shocks form when the SN shock reaches the outer edge of the He core very soon after the start of the simulation. The reverse shock then propagates slowly inward (in mass space) and will eventually reach the innermost ejecta layers after about 10 d (as indicated in the V1D simulation; not shown here). The reverse shock plows through the extended envelope more swiftly but it undergoes a strong deceleration as because of the relatively flat density profile of the massive extended H-rich envelope. Over the course of the simulation, the ejecta density drops from $10^{-4}$ to $10^{-7}$ g cm-3, primarily through the expansion of the dense inner regions over a timescale of days. In contrast, the outer regions, located at large distances from the explosion site, are shocked last and undergo limited expansion. A high-density plateau forms around the peak density so that most of the shocked-envelope mass is located between 500 and 900 $R_{\odot}$ at shock emergence. A small and decreasing mass is present below the reverse shock owing to the large progenitor envelope mass relative to the He-core mass.
The velocity profile mirrors that of the density. From the time of the explosion, the peak velocity decreases from 7500 to 3000 km s-1. Over time, the shock front remains steep while propagating through the envelope. The main difference between the two codes is the prediction of a temperature drop interior to the reverse shock. This feature is also present in MESA but absent in V1D because of the use of a temperature (or internal energy) floor in the EoS (for a similar pattern, see Fig. 28 of Paxton et al. 2015). There is a good agreement between the two codes for the temperature in the shocked regions, which are radiation dominated (i.e., $P_{\rm r}\gg P_{\rm g}$).
@@
Figure 11: Multiepoch meridional slices for the 3-D RSG-star explosion simulation with HERACLES++. The panels show the radial velocity (a), the normalized density $\rho/\langle\rho\rangle$ (b), the polar velocity $u_{\theta}$ (c), and the He mass fraction $X_{\rm He}$ (d). Within each panel, each quadrant corresponds to a cut at $\phi=\pi/2$ and progresses clockwise from 1, to 4, 8 and $12\times 10^{4}$ s. (See §5.5.2 for discussion.)
@@ 5.4 Stability analysis
Before discussing the 3-D simulations of our explosion model, we conduct an approximate local stability analysis using the method of Benz & Thielemann (1990) – see also Mueller et al. (1991). It estimates the growth rate of the Rayleigh-Taylor instability in the compressible case as:
@@
Similarly, the total time-integrated growth at a given mass is given by:
@@
Figure 10 shows the integrated growth rate as a function of Lagrangian mass at selected epochs (in the incompressible case, $\xi/\xi_{0}$ is qualitatively similar but peaks at smaller values). As expected, the most unstable region is located at the interface between the highly bound He-core and the loosely bound H-rich envelope, at $\sim$ 4 $M_{\odot}$, and thus where there is a density cliff. The small peak at $\sim$ 1.5 $M_{\odot}$ is an artifact of the initial ejecta density structure reflecting the regions of the original thermal bomb. This analysis indicates that the Rayleigh-Taylor instability should develop most strongly at the H-He interface in 3-D simulations, as previously found (see, for example, Mueller et al. 1991).
5.5 Results from the 3-D simulation

### v2

@@ 4.3 Three-dimensional off-centered Sedov blast wave test
The test of the Sedov blast wave (Sedov, 1946) is useful to describe the energy-conserving phase of a SN remnant. We performed this test using a grid in spherical geometry in order to test the precision of our numerical scheme for a non-Cartesian grid. The problem consists of a stationary background medium with a uniform density, $\rho_{0}=1$, and a low pressure, $P_{0}=0.01$. For our test, the domain was a spherical wedge with $r\in[0.1,1.3]$ and $(\theta,\phi)\in[\pi/4,3\pi/4]$. The blast was created with a total energy peak of $E_{1}=1$ at the position $\vec{r_{0}}=(r_{0},\theta_{0},\phi_{0})=(0.7,\pi/2,\pi/2)$. The boundary conditions are transmissive in every direction, although the simulation is stopped before the blast wave reaches any boundary. The spherical polar grid counts 256 zones in all three directions.
The distance between the shock and $\vec{r_{0}}$ is given analytically by
$d_{\rm sh}(t)=\xi_{0}\Big{(}\frac{E_{1}\,t^{2}}{\rho_{0}}\Big{)}^{1/5},$ (25)
@@
Figure 5 shows a meridional slice of the density at $t=0.08$ together with the analytical solution given by Eq. 25. Other slices through the 3D domain yield a similar pattern. Hence, the simulation is in good agreement with the analytical prediction, apart from a small offset in the direction where the shock is aligned with the grid. In such locations, the numerical diffusion is lower, causing a notorious problem for numerical schemes (Elling, 2009).
Figure 5: Meridional slice of the density for the 3D off-centered Sedov blast wave test. The time is 0.08 and the cut is done along $\phi=\pi/2$. The analytical prediction given by Eq. 25 is shown as a black line. (See Sect. 4.3 for discussion.)
@@
Figure 6: Initial conditions for our HERACLES++ simulations. We show the density (top left), temperature (top right), velocity (bottom left), and the mass fraction of H, He, O, Si, and 56Ni, as a function of Lagrangian mass. The time is 500 s after the explosion trigger.
@@ 5.1 Preamble
We now turn to the more astrophysics-oriented context of a spherical explosion occurring in a RSG star. This initial work is used here for the benchmarking of the code but it fits within a larger project of simulating the various instabilities taking place in the extended envelopes of RSG stars before, during, and after their explosion. These so-called type II SNe are the most frequently encountered type of SN from massive star explosions (see, e.g., Perley et al. 2020). Furthermore, detailed NLTE radiative-transfer calculations have demonstrated that both small and large-scale structures of SN ejecta have a significant impact on SN light curves and spectra (Dessart et al., 2018; Dessart & Audit, 2019; Ergon & Fransson, 2022) so progress in that sector is much needed in order to improve the physical consistency of radiative-transfer and radiation-hydrodynamics simulations and the robustness of their inferences.
For this work, we used the initial ejecta conditions of Dessart et al. (2024), which are based on a subset of evolutionary calculations for binary massive stars performed by Ercolino et al. (2024). Binarity, which occurs with a high frequency in massive stars (Sana et al., 2012), can lead to a wide range of H-rich envelope masses, $M$(H-env), in the pre-SN star because of mass transfer through Roche-lobe overflow. However, as long as a residual H-rich envelope remains, the He-core mass as well as the surface radii are largely unaffected by this mass loss, so that these objects should explode with similar properties (explosion energy and nucleosynthesis), merely modulated by the differing $M$(H-env). Dessart et al. (2024) indeed showed that through variations in $M$(H-env) alone, the full diversity of type II SN light curves may be reproduced. Chemical mixing and clumping may, however, differ widely between these ejecta due to the differences in their progenitor-envelope properties, with strong implications for their observables. Investigating this diversity will be the subject of a forthcoming study.
@@
In practice, we started our simulations from the output of the explosion simulation carried out with the 1D, Lagrangian radiation-hydrodynamics code V1D (Livne, 1993; Dessart et al., 2010a, b), at a time of 500 s relative to the onset of the explosion trigger. The total energy injected in the star for the explosion was designed to yield a kinetic energy at infinity of $1.1\times 10^{51}$ erg so by this time of 500 s, the shock has already reached the outer layers of the He core [^3] For details about the preSN evolution and explosion simulation, we refer the reader to Dessart et al. (2024) and Ercolino et al. (2024).. Using HERACLES++, we first evolved this structure in 1D (Sect. 5.3) and then in 3D (Sect. 5.5) in order to study the development of Rayleigh-Taylor instabilities and their impact on chemical mixing and clumping.
Figure 7: Variation of the radial grid spacing, $dr$, as the function of the radius, $r$. The grid used in the simulation with HERACLES++ is regular until 10 $R_{\odot}$ and logarithmic beyond.
Figure 8: Ejecta properties for our RSG star explosion model as computed by HERACLES++ (solid) and V1D (dashed). For representative epochs between 2.8 hr and 1.6 d, we show the density (top row), the temperature (middle row), and the velocity (bottom row) vs. the Lagrangian mass (left column) and radius (right column). (See Sect. 5.3 for discussion.)
@@ 5.2 Initial conditions and numerical setup
The ejecta properties used as initial conditions for the HERACLES++ simulations are shown in Fig. 6. The shock is located at $\sim$ 4.1 $M_{\odot}$ with a peak velocity around 12 000 km s-1. Beyond the shock, the velocity is essentially zero and $\rho$ and $T$ slowly decrease until the progenitor surface at 10.84 $M_{\odot}$. Figure 6 also shows the mass fraction of the main elements (i.e., H, He, O, Si, and ${}^{56}\rm{N}$i). The position of each layer in the star reflects previous episodes of nuclear burning or mixing, as well as the ashes of explosive nucleosynthesis in the innermost ejecta layers. The extended H-rich envelope is essentially unchanged from its original composition (i.e., corresponding here to solar metallicity), whereas the He-core layers exhibit a stratification of ashes from H, He, C, O, Ne, and Si burning. The explosion energy at 500 s is primarily stored as kinetic energy, with around $9\times 10^{50}$ erg.
@@
Since our interest is to investigate the impact of fluid instabilities on the distribution of elements, we introduced five passive scalars to track the material originally present in the H-rich envelope and in the He-, O-, Si-, or ${}^{56}\rm{N}$i-rich shells. In the following sections, when referring to the distribution of O, we really mean the distribution of the material in the O-rich shell (in practice, this includes the O/Si, O/Ne/Mg, and the O/C shells).
@@
Figure 9: Evolution of the total energy, $E_{\rm tot}$, the kinetic energy, $E_{\rm kin}$, the radiative energy, $E_{\rm rad}$, and the internal (gas) energy, $E_{\rm gas}$, during our HERACLES++ simulation from 500 s until shock breakout. At shock breakout, $E_{\rm kin}\sim E_{\rm rad}$.
@@
At the start of the HERACLES++ simulation, we remapped the V1D input (which is on a Lagrangian grid) onto the HERACLES++ grid, which we chose to be a hybrid linear-log grid. It has a relatively high resolution out to the shock location (in order to preserve the initial sharpness of the shock and the total initial energy of the explosion) and a decreasing resolution beyond. In Sect. 5.3, we constructed a regular radial grid with 10000 points to produce a grid that presents no major resolution offset with the V1D grid, which is Lagrangian and thus in mass space (i.e., the goal in this test is to confront the numerical solution of the hydrodynamical equations between the two codes, rather than assess the influence of resolution or numerical diffusion etc). In Sect. 5.5, we created a radial grid with a total of 1024 points, with 200 points for the regular part out to 10 $R_{\odot}$ (corresponding to a Lagrangian mass of $\sim$ 4.2 $M_{\odot}$). Figure 7 illustrates the variation in the radial increment, $dr$, with radius. In Appendix A, we consider additional radial resolutions (i.e., $n_{r}$ of 256, 512, and 768) as part of a resolution study.
@@
We performed hydrodynamical simulations with HERACLES++, using spherical coordinates, first in 1D (Sect. 5.3) and then in 3D (Sect. 5.5). We used the perfect gas+radiation EoS (Eq. 6b; we used $\gamma=5/3$ and $\mu=0.62$ for the ideal-gas part) and introduced ${}^{56}\rm{N}$i-decay heating (Eq. 8; this was done for consistency with V1D but it plays little role here since we stopped the simulation at about 1.5 d, which is only a quarter of the ${}^{56}\rm{N}$i lifetime). A point mass gravity (Eq. 3) was used to account for the presence of the compact remnant. We used a routine to track precisely the internal energy (Sect. 2.5). The boundary conditions were reflexive in the inner radius and transmissive at the outer radius. The slope limiter used was Minmod and the Riemann solver was HLLC. For consistency, this general setup was the same for all similar simulations.
@@
Figure 10: Estimated growth rate of the R-T instability in our RSG explosion simulation, shown at multiple epochs and vs. Lagrangian mass. (See Sect. 5.4 for discussion.)
@@ 5.3 Results in spherical symmetry and comparison with V1D
In this section, we first present a comparison of the results between V1D and HERACLES++, and thus in spherical symmetry. Figure 8 shows the evolution of $\rho$, $T$, and $u$ over time as the shock propagates through the progenitor envelope and until its emergence at $R_{\star}$ at $\sim$ 1.6 d. [^4] For this figure and to ensure the Eulerian grid is as fine in the inner regions as the Lagrangian grid in V1D, the HERACLES++ simulation employs 10 000 radial zones. We show the variations in these quantities as a function of both the Lagrangian mass and the radius. There is a good agreement between the results of the two codes.
A pair of reverse and forward shocks forms when the SN shock reaches the outer edge of the He core very soon after the start of the simulation. The reverse shock then propagates slowly inward (in mass space) and will eventually reach the innermost ejecta layers after about 10 d (as is indicated in the V1D simulation; not shown here). The reverse shock plows through the extended envelope more swiftly but it undergoes a strong deceleration because of the relatively flat density profile of the massive extended H-rich envelope. Over the course of the simulation, the ejecta density drops from $10^{-4}$ to $10^{-7}$ g cm-3, primarily through the expansion of the dense inner regions over a timescale of days. In contrast, the outer regions, located at large distances from the explosion site, are shocked last and undergo limited expansion. A high-density plateau forms around the peak density so that most of the shocked-envelope mass is located between 500 and 900 $R_{\odot}$ at shock emergence. A small and decreasing mass is present below the reverse shock owing to the large progenitor envelope mass relative to the He-core mass.
The velocity profile mirrors that of the density. From the time of the explosion, the peak velocity decreases from 7500 to 3000 km s-1. Over time, the shock front remains steep while propagating through the envelope. The main difference between the two codes is the prediction of a temperature drop interior to the reverse shock. This feature is also present in MESA but absent in V1D because of the use of a temperature (or internal energy) floor in the EoS (for a similar pattern, see Fig. 28 of Paxton et al. 2015). There is a good agreement between the two codes for the temperature in the shocked regions, which are radiation-dominated (i.e., $P_{\rm r}\gg P_{\rm g}$).
@@
Figure 11: Multiepoch meridional slices for the 3D RSG-star explosion simulation with HERACLES++. The panels show the radial velocity (a), the normalized density, $\rho/\langle\rho\rangle$ (b), the polar velocity, $u_{\theta}$ (c), and the He mass fraction, $X_{\rm He}$ (d). Within each panel, each quadrant corresponds to a cut at $\phi=\pi/2$ and progresses clockwise from 1, to 4, 8, and $12\times 10^{4}$ s. (See Sect. 5.5.2 for discussion.)
@@ 5.4 Stability analysis
Before discussing the 3D simulations of our explosion model, we conduct an approximate local stability analysis using the method of Benz & Thielemann (1990) (see also Mueller et al. (1991)). It estimates the growth rate of the Rayleigh-Taylor instability in the compressible case as:
@@
Similarly, the total time-integrated growth at a given mass is given by
@@
Figure 10 shows the integrated growth rate as a function of the Lagrangian mass at selected epochs (in the incompressible case, $\xi/\xi_{0}$ is qualitatively similar but peaks at smaller values). As is expected, the most unstable region is located at the interface between the highly bound He-core and the loosely bound H-rich envelope, at $\sim$ 4 $M_{\odot}$, and thus where there is a density cliff. The small peak at $\sim$ 1.5 $M_{\odot}$ is an artifact of the initial ejecta density structure reflecting the regions of the original thermal bomb. This analysis indicates that the Rayleigh-Taylor instability should develop most strongly at the H-He interface in 3D simulations, as was previously found (see, for example, Mueller et al. 1991).
5.5 Results from the 3D simulation